# Benchmarking XAI Explanations with Human-Aligned Evaluations

### v1

@@ 4 Developing a Metric for Perceptual Evaluation
Our human preference dataset contains evaluations of the fidelity, complexity, objectivity, and robustness of each evaluation. These scores were painstakingly attributed by human annotators. To provide a tool for measuring human assessment of XAI techniques, we introduce a model, the PASTA-metric, which simulates human evaluation. The global pipeline is illustrated in Figure 4. More precisely, the PASTA-metric is composed of an embedding network (Section 4.1), that processes both CBM outputs or saliency maps, and a scoring network (Section 4.2), that computes scores from the embeddings. Further details on the implementation are given in Appendix C.1. Using the data collected in Section 3, the PASTA-metric aims at predicting the human scores for questions 1 to 6, aggregated by majority voting.
Figure 4: Pipeline of the PASTA-metric. Initially, embeddings are computed based on the explanations. Then a scoring network, trained on labels provided by the Benchmark Dataset, generates a final score.
4.1 Computation of embeddings
Drawing inspiration from recent literature in automated scoring (Yang et al., 2020; Wang et al., 2022b), we use a foundation model to generate embeddings. Given its multimodal capabilities, we select CLIP (Yan et al., 2023) as the embedding model. This choice allows for a unified integration of both concept-based explanations, which can be transformed into text, and saliency map-based explanations, which can be projected into the same embedding space.
Let us denote by ${\bm{x}}_{i}\in\mathbb{R}^{H\times W\times 3}$ the $i$-th test image of height $H$ and width $W$ in the dataset and by ${\bm{e}}_{i}^{saliency}\in\mathbb{R}^{H\times W}$ any explanation produced by a saliency-based XAI method for this image. We also note the CLIP image encoder as $\text{CLIP}_{\text{image}}$. For saliency explanations, the resulting embedding based on a saliency map can be obtained using the following formula:
$\phi_{\text{image}}({\bm{e}}_{i}^{saliency})=\text{CLIP}_{\text{image}}(Heatmap({\bm{x}}_{i},{\bm{e}}_{i}^{saliency})),$ (1)
where $Heatmap$ is the process generating the saliency related heatmap to the image. More details are given in Appendix C.3.1.
For CBMs, let us denote with ${\bm{e}}_{i}^{CBM}\in\mathbb{R}^{K}$ any explanation produced by a saliency-based XAI method for this image ${\bm{x}}_{i}$, where $K$ is the length of the concept set. These attributions are converted into a sentence, which is then embedded. Let $\text{CLIP}_{\text{text}}$ denote the CLIP text encoder and $Sentence$ the process of converting the CBM explanation into text, more details about which $Sentence$ are given in Appendix C.3.1. The resulting embedding is:
$\phi_{\text{text}}({\bm{e}}_{i}^{CBM})=\text{CLIP}_{\text{text}}(Sentence({\bm{e}}_{i}^{CBM}))\,.$ (2)
4.2 Scoring network
@@
Once the embeddings are computed, a scoring network composed of a linear layer is used to predict scores. Inspired by Automated Essay Scoring  (Yang et al., 2020; Wang et al., 2022b), we use a loss $L$ that combines a similarity loss $L_{s}$, a mean squared error (MSE) loss $L_{mse}$, and a ranking loss $L_{r}$. From a set of ground truth scores obtained from majority voting $\{m_{k}\}_{k\in[0,N_{s}]}$ and the predictions given by the scoring network $\{\hat{m}_{k}\}_{k\in[0,N_{s}]}$, the resulting loss is defined as:
$L=\alpha L_{s}+\beta L_{mse}+\gamma L_{r},$ (3)
@@
where $\alpha$, $\beta$, and $\gamma$ are hyperparameters controlling the relative importance of each component. Details about the different losses are given in Appendix C.3
4.3 Classifier results
In all experiments, we employed the Adam optimizer with a batch size of 128, training for 500 epochs at a learning rate of 0.001. Additionally, we configured the parameters as follows: $\alpha$ was set to 1, $\beta$ to 0.01, and $\gamma$ to 0.1. We split our dataset in $1540$ training samples, $330$ validation samples, and $330$ testing samples. The experiments were conducted using a V100-16GB GPU. The training and inference times are summarized in Table 2. It is important to note that, for both inference and testing, the majority of the computational time is dedicated to precomputing CLIP embeddings, as the scoring network itself is relatively lightweight and requires minimal computational resources.
Table 2: Training and Inference Times on GPU and CPU.
Device Training Time (s) Inference Time (s)
GPU 214.18 8.89
CPU 1308.84 57.52
For questions 1 to 6 in Section 3, we calculated the Mean Square Error (MSE), Quadratic Weighted Kappa (QWK), and Spearman Correlation Coefficient (SCC) between the predicted and ground truth labels on the test set. The results are presented in Table 3. Additional details about these computations can be found in Appendix C.1.2. For comparison, we also include the scores from a linear regression model and report the inter-annotator agreement values, which correspond to the metrics computed between a randomly selected annotator’s score and the mode. Additional experiments, including ablation studies, are presented in Appendix C.2.
@@ 4.3 Classifier results
Table 3: Mean Square Error, Quadratic Weighted Kappa and Spearman Correlation Coefficient for each question. Each value is the average result on 5 runs with the standard deviation. Human here stands for the inter-annotator agreement.
Regressor Quest. MSE QWK SCC
CLIP+Linear Regression Q1 1.48 $\pm$ 0.13 0.37 $\pm$ 0.04 0.19 $\pm$ 0.19
PASTA Q1 0.85 $\pm$ 0.06 0.55 $\pm$ 0.04 0.28 $\pm$ 0.28
Human Q1 0.53 $\pm$ 0.03 0.73 $\pm$ 0.03 0.37 $\pm$ 0.37
CLIP+Linear Regression Q2 1.45 $\pm$ 0.14 0.36 $\pm$ 0.04 0.19 $\pm$ 0.19
PASTA Q2 0.86 $\pm$ 0.06 0.54 $\pm$ 0.03 0.28 $\pm$ 0.28
Human Q2 0.51 $\pm$ 0.05 0.74 $\pm$ 0.02 0.38 $\pm$ 0.38
CLIP+Linear Regression Q3 1.55 $\pm$ 0.14 0.33 $\pm$ 0.03 0.17 $\pm$ 0.17
PASTA Q3 0.96 $\pm$ 0.10 0.51 $\pm$ 0.05 0.26 $\pm$ 0.26
Human Q3 0.74 $\pm$ 0.03 0.63 $\pm$ 0.02 0.33 $\pm$ 0.33
CLIP+Linear Regression Q4 1.46 $\pm$ 0.10 0.34 $\pm$ 0.05 0.18 $\pm$ 0.18
PASTA Q4 0.90 $\pm$ 0.05 0.52 $\pm$ 0.03 0.27 $\pm$ 0.27
Human Q4 0.72 $\pm$ 0.02 0.62 $\pm$ 0.02 0.33 $\pm$ 0.33
CLIP+Linear Regression Q5 2.82 $\pm$ 0.22 0.14 $\pm$ 0.05 0.10 $\pm$ 0.08
PASTA Q5 1.86 $\pm$ 0.16 0.29 $\pm$ 0.08 0.15 $\pm$ 0.16
Human Q5 1.00 $\pm$ 0.06 0.65 $\pm$ 0.03 0.34 $\pm$ 0.34
CLIP+Linear Regression Q6 1.02 $\pm$ 0.06 0.32 $\pm$ 0.05 0.15 $\pm$ 0.16
PASTA Q6 0.67 $\pm$ 0.09 0.51 $\pm$ 0.08 0.25 $\pm$ 0.26
Human Q6 0.52 $\pm$ 0.03 0.59 $\pm$ 0.02 0.29 $\pm$ 0.29
4.4 Generalization capabilities of the PASTA-metric
In the main study, to constitute training, validation, and test sets, we shuffled all the samples without considering the XAI technique or the dataset they belong to. This means that the same images with different XAI techniques, or the same XAI techniques applied to different images, possibly from the same dataset, could be found in the training and test set. In this section, we investigate the impact of shuffling solely the image indices and the explanation indices. By doing so, we ensure that samples from the same image (resp. same XAI technique) cannot be in two different splits. This will help us investigate the generalization capabilities of the model in two distinct ways: can it generalize to new XAI techniques, and to new images? The results of the two setups for Q1 are shown in Table 4. The results indicate a decrease of 0.04 in QWK when shuffling across XAI techniques and a more significant drop of 0.07 when shuffling across image IDs. This opens a discussion on the potential for applying the PASTA-metric to other image datasets. Regarding the generalization to new XAI methods, the relatively moderate drop in performance supports the feasibility of testing our metric on novel XAI techniques.
Table 4: Summary of results for different restrictions applied during dataset splitting. The label No indicates no restrictions, Img id denotes that the same image indices $i$ are maintained across splits, and XAI id indicates that the same explanation indices $j$ are preserved in different splits. Each value is the average result on 5 runs with the standard deviation.
Restriction split MSE QWK SCC
No 0.85 $\pm$ 0.06 0.55 $\pm$ 0.04 0.28 $\pm$ 0.28
xai_id 0.96 $\pm$ 0.10 0.51 $\pm$ 0.05 0.26 $\pm$ 0.26
img_id 1.06 $\pm$ 0.05 0.48 $\pm$ 0.05 0.25 $\pm$ 0.25
5 Conclusions
In this paper, we introduce PASTA, a novel perceptual assessment system designed to benchmark explainable AI (XAI) techniques in a human-centric manner. We integrate four diverse datasets — COCO, Pascal Parts, Cats Dogs Cars, and Monumai — to form a large-scale benchmark dataset for XAI, and used it for an assessment of XAI explanations by human annotators. We also develop an automated evaluation metric that mimics human preferences based on a comprehensive database of human evaluations. This framework offers a scalable and reliable way to compare different XAI methods, facilitating robust evaluations across modalities previously unaddressed.
Our findings demonstrate a clear preference for saliency-based explanations, particularly techniques such as LIME and SHAP, which align well with human intuition. These results affirm the scalability and reliability of our perceptual metric, which provides consistency with human assessment while automating much of the evaluation process.
However, there are limitations to our approach. The current study focuses on a fixed set of datasets and XAI techniques. Human evaluations can be influenced by subjective factors that may affect the consistency of results. Furthermore, annotators can inadvertently introduce biases. Perceptual metrics like the ones proposed here are therefore not intended to serve as absolute measures of XAI performance. Rather, we consider them simply as complementary to other XAI metrics.

### v2

@@ Introduction
To provide a tool for measuring human assessment of XAI techniques, we introduce the PASTA-score, which simulates a human evaluation. The global pipeline is illustrated in Figure 6. More precisely, the PASTA-score is composed of an embedding network, that processes both CBM outputs or saliency maps, and a scoring network, that computes scores from the embeddings. Using the data collected in the PASTA-dataset, the PASTA-score aims at predicting the human scores for questions Q1 to Q6 for new explanations, playing the role of an automated benchmark.
Figure 6: Functioning of the PASTA-score. First, we extract the embeddings for each explanation using a frozen image encoder of a VLM. Then, we employ a scoring network trained on the labels provided by the PASTA-dataset to generate a final score.
Computation of the embeddings
Drawing inspiration from recent literature in automated essay scoring (Yang et al. 2020; Wang et al. 2022), which encounters similar challenges due to working with a dataset of thousands of samples (21,110 samples per question, precisely) and requiring a DNN to automatically learn the score, we opt for a foundation model that we will fine-tune using a multi-linear perceptron. However, unlike automated essay scoring, we deal with both image and textual inputs, making the use of a Vision Language Model (VLM) mandatory. We tested multiple candidates, such as CLIP (Yan et al. 2023), BLIP (Li et al. 2022a) and LLaVa (Liu et al. 2024b). This choice allows for a unified integration of both concept-based explanations, which can be transformed into text, and saliency map-based explanations, which can be projected into the same embedding space. Let ${\bm{x}}_{i}\in\mathbb{R}^{H\times W\times 3}$ be the $i$-th test image with height $H$ and width $W$, and let ${\bm{e}}_{i}^{saliency}\in\mathbb{R}^{H\times W}$ be a saliency-based explanation for this image. We denote the image encoder of a Vision-Language Model (VLM) as $\text{VLM}_{\text{image}}$. To embed a saliency explanation, we apply the encoder to the image overlaid with its heatmap:
$\phi_{\text{image}}({\bm{e}}_{i}^{saliency})=\text{VLM}_{\text{image}}(Heatmap({\bm{x}}_{i},{\bm{e}}_{i}^{saliency})),$ (1)
where $Heatmap$ generates the visual overlay of the explanation on the image.
For concept-based methods (CBMs), let ${\bm{e}}_{i}^{CBM}\in\mathbb{R}^{K}$ be the explanation vector, where $K$ is the number of concepts. This vector is turned into a sentence using a text template, and then encoded with the VLM’s text encoder $\text{VLM}_{\text{text}}$:
$\phi_{\text{text}}({\bm{e}}_{i}^{CBM})=\text{VLM}_{\text{text}}(Sentence({\bm{e}}_{i}^{CBM})).$ (2)
Since the PASTA-score is compatible with different VLMs and does not rely on a specific one, we evaluate it using several VLMs: CLIP (Liu et al. 2024b), SIGLIP (Zhai et al. 2023), EVA (Sun et al. 2023), and BLIP (Li et al. 2022a). This results in multiple variants of our metric: $\text{PASTA-score}^{\text{CLIP}}$, $\text{PASTA-score}^{\text{SIGLIP}}$, $\text{PASTA-score}^{\text{EVA}}$, and $\text{PASTA-score}^{\text{BLIP}}$.
To support our design choices, Appendix D.2 presents extensive ablations on various factors: the impact of textual templates, the number of concepts $K$, the way saliency maps are visualized, and whether label information is included. We also explore alternative versions of Equations 1 and 2, and how these choices affect the final score.
Scoring network
@@
Once the embeddings are computed, the label information is concatenated to the embedding, and a scoring network composed of a multi-layer perceptron is used to predict scores. Inspired by Automated Essay Scoring  (Yang et al. 2020; Wang et al. 2022), we use a loss $L$ that combines a similarity loss $L_{s}$, a mean squared error (MSE) loss $L_{mse}$, and a ranking loss $L_{r}$. From a set of ground truth scores obtained from majority voting $\{m_{k}\}_{k\in[0,N_{s}]}$ and the predictions given by the scoring network $\{\hat{m}_{k}\}_{k\in[0,N_{s}]}$, the resulting loss is defined as:
$L(m_{k},\hat{m}_{k})=\alpha L_{s}(m_{k},\hat{m}_{k})+\beta L_{mse}(m_{k},\hat{m}_{k})$
$+\gamma L_{r}(m_{k},\hat{m}_{k}),$ (3)
@@
where $\alpha$, $\beta$, and $\gamma$ are hyperparameters controlling the relative importance of each component. Formulas about the different losses are given in Appendix D.1. Since the PASTA dataset provides 5 ground-truth votes per inference, we explored different aggregation strategies. To mitigate the phenomenon of high non-consensus, the mode was selected as the final choice.
Classifier results
Note that in the PASTA-dataset each sample corresponds to a triplet (input image, explanation, human ratings). The same image thus appears multiple times for different XAI methods, and the same XAI method appears multiple times for different images. To guarantee that no leakage occurs between train-test splits, we design them to ensure that the same image, or the same XAI method, does not appear in different splits. Images and XAI methods included in the training splits are randomly chosen based on the run’s random seed. For Q1 to Q6, we calculate the Mean Square Error (MSE), Quadratic Weighted Kappa (QWK), and Spearman Correlation Coefficient (SCC) between the predicted and ground truth labels on the test set. The results are presented in Table 1, where we also ablate different choices of embedding methods. We also report the inter-annotator agreement values, which correspond to the expected deviation of the metrics between a randomly selected annotator’s score and the mode. Our network best replicates answers to Q1 and Q2, with similar performance across $\text{PASTA-score}^{\text{CLIP}}$ and $\text{PASTA-score}^{\text{SIGLIP}}$. This is likely due to greater rating diversity and stronger agreement between annotators, which supports more stable training. In contrast, Q3 to Q5 shows lower agreement, and Q5–Q6 has less diverse ratings. While the MSE stays similar, it becomes harder to learn the ranking patterns, likely due to the more subjective nature of these questions and the added uncertainty from image perturbations in Q5 and Q6.
@@
Table 1: Mean Square Error (MSE), Quadratic Weighted Kappa (QWK), and Spearman Correlation Coefficient (SCC) for each question. Each value is the average of 5 runs with standard deviation. Human refers to inter-annotator agreement.
Metric Model Q1 Q2 Q3 Q4 Q5 Q6
MSE $\downarrow$ $\text{PASTA-score}^{\text{CLIP}}$ 0.990 $\pm$ 0.104 0.993 $\pm$ 0.096 2.111 $\pm$ 2.529 0.811 $\pm$ 0.095 1.476 $\pm$ 0.183 0.752 $\pm$ 0.127
$\text{PASTA-score}^{\text{SIGLIP}}$ 0.989 $\pm$ 0.113 1.009 $\pm$ 0.125 0.842 $\pm$ 0.094 0.840 $\pm$ 0.106 1.396 $\pm$ 0.177 0.739 $\pm$ 0.140
$\text{PASTA-score}^{\text{BLIP}}$ 3.297 $\pm$ 1.840 3.287 $\pm$ 1.835 5.938 $\pm$ 2.542 4.642 $\pm$ 3.135 3.005 $\pm$ 1.385 10.710 $\pm$ 4.943
$\text{PASTA-score}^{\text{EVA}}$ 1.666 $\pm$ 1.215 1.747 $\pm$ 1.174 3.355 $\pm$ 3.099 2.097 $\pm$ 2.608 1.767 $\pm$ 0.568 3.324 $\pm$ 5.091
Human 0.415 $\pm$ 0.037 0.429 $\pm$ 0.049 0.562 $\pm$ 0.104 0.478 $\pm$ 0.080 0.509 $\pm$ 0.102 0.299 $\pm$ 0.051
QWK $\uparrow$ $\text{PASTA-score}^{\text{CLIP}}$ 0.450 $\pm$ 0.066 0.452 $\pm$ 0.063 0.199 $\pm$ 0.040 0.216 $\pm$ 0.052 0.165 $\pm$ 0.060 0.159 $\pm$ 0.031
$\text{PASTA-score}^{\text{SIGLIP}}$ 0.471 $\pm$ 0.055 0.459 $\pm$ 0.056 0.237 $\pm$ 0.052 0.219 $\pm$ 0.035 0.177 $\pm$ 0.061 0.165 $\pm$ 0.018
$\text{PASTA-score}^{\text{BLIP}}$ 0.328 $\pm$ 0.023 0.340 $\pm$ 0.020 0.181 $\pm$ 0.003 0.173 $\pm$ 0.017 0.081 $\pm$ 0.081 0.159 $\pm$ 0.011
$\text{PASTA-score}^{\text{EVA}}$ 0.462 $\pm$ 0.050 0.457 $\pm$ 0.054 0.160 $\pm$ 0.099 0.230 $\pm$ 0.018 0.163 $\pm$ 0.029 0.185 $\pm$ 0.049
Human 0.849 $\pm$ 0.013 0.845 $\pm$ 0.017 0.731 $\pm$ 0.050 0.748 $\pm$ 0.041 0.848 $\pm$ 0.029 0.796 $\pm$ 0.048
SCC $\uparrow$ $\text{PASTA-score}^{\text{CLIP}}$ 0.484 $\pm$ 0.064 0.484 $\pm$ 0.062 0.213 $\pm$ 0.040 0.230 $\pm$ 0.050 0.197 $\pm$ 0.073 0.193 $\pm$ 0.030
$\text{PASTA-score}^{\text{SIGLIP}}$ 0.501 $\pm$ 0.052 0.490 $\pm$ 0.057 0.247 $\pm$ 0.048 0.223 $\pm$ 0.029 0.213 $\pm$ 0.071 0.196 $\pm$ 0.020
$\text{PASTA-score}^{\text{BLIP}}$ 0.397 $\pm$ 0.036 0.411 $\pm$ 0.033 0.207 $\pm$ 0.065 0.194 $\pm$ 0.014 0.088 $\pm$ 0.104 0.218 $\pm$ 0.012
$\text{PASTA-score}^{\text{EVA}}$ 0.484 $\pm$ 0.046 0.474 $\pm$ 0.048 0.150 $\pm$ 0.155 0.247 $\pm$ 0.015 0.216 $\pm$ 0.035 0.220 $\pm$ 0.057
Human 0.844 $\pm$ 0.017 0.839 $\pm$ 0.019 0.722 $\pm$ 0.045 0.742 $\pm$ 0.038 0.850 $\pm$ 0.023 0.789 $\pm$ 0.039
Alignment with Established Benchmarks
In this subsection, we aim to evaluate whether the PASTA-score aligns with the findings of existing studies that compare human assessments of XAI methods. To accomplish this, we use the benchmark dataset established by Yang et al. (2022), which offers comparative evaluations of GradCAM (Selvaraju et al. 2017), RISE (Petsiuk, Das, and Saenko 2018) (at the image level), and Guided Backpropagation (Springenberg et al. 2014). These quantitative evaluations are presented in the form of Mean Absolute Error (MAE) between generated saliency maps and those produced by a Human Saliency Imitator (HSI). Given that this benchmark aims to explore human expectations, we compare the results obtained by the authors (MAE) with the PASTA-score related to Q1.
Table 2: Comparison of Rankings Based on HSI and PASTA-Scores. The results illustrate that the PASTA-score exhibits a correlation with the Human Saliency Imitator presented by (Yang et al. 2022).
Method MAE Score $\downarrow$ PASTA-score $\uparrow$
RISE 0.442 3.113
GradCAM 0.703 3.150
Guided Backpropagation 0.890 1.846
The results are summarized in Table 2, and indicate that despite RISE and Guided Backpropagation not being included during the training of the PASTA metric, the obtained rankings remain consistent. This provides empirical evidence supporting the generalization of the PASTA-score to unseen XAI techniques.
Applications
In this section, we explore three different applications using the PASTA-score as a replacement for human feedback, which would be difficult or too costly to run at scale without automation We use PASTA-score to guide XAI methods toward better interpretability (in the first and third applications) and to analyze how model size affects interpretability (in the second application). All experiments use the PASTA-score model trained on Q1 for consistency.
Mixture of XAI methods
Our first application is to dynamically select the explainer giving the explanation that best matches human judgments for each specific image, using a mixture of XAI methods. In our experiments, we fix the classifier to be a ResNet50, and we select the explanation with the highest PASTA-score for each image. The distribution of selected XAI methods is shown in Table 7.
Figure 7: Distribution of methods selected by our mixture of XAI techniques. The x-axis denotes the number of images in the automated benchmark for which the respective XAI method attained the highest performance.
The results indicate a substantial diversity in the methods employed, with FullGrad emerging as the most frequently used, selected nearly half of the time. This trend is reflective of user ratings within the PASTA-dataset, where FullGrad is identified as providing the most effective explanations according to annotators. In terms of faithfulness, the computation of the average faithfulness correlation across explanations selected by our PASTA-score yields a relatively stable value, with a slight improvement compared to the value obtained by averaging over every explainer (0.0627 for our selection versus 0.0579 for the average over every explainer). This confirms that it is possible to use PASTA to enhance the interpretability of explanations without compromising their faithfulness.
Backbone size influence on the understanding of explanation
We use the PASTA-score to investigate whether model size influences the human perception of explanations, and how this relates to XAI methods. To this end, we compute the average PASTA-score within an identical experimental framework, varying only the size of the backbone model. Specifically, we employ CLIP as the classifier and select backbones from among its ViT-B-16, ViT-L-14, ViT-H-14, and ViT-g-14 variants. The results of this analysis are presented in Figure 8.
(a) Comparison of various XAI methods across different ViT models.
(b) ViT-B-16; PASTA-score of the explanation: 3.12
(c) ViT-g-14; PASTA-score of the explanation: 2.75
Figure 8: Impact of classifier backbone size on the perceived interpretability of image explanations. A notable decrease in the PASTA-score is observed as the model size increases (left). Examination of image samples suggests that artifacts present in the background are likely responsible for this decline.
Our results show that for activation map-based XAI methods, performance metrics drop as the model size increases. This decline is particularly pronounced when transitioning from the ViT-B to the ViT-L architecture. Several hypotheses may account for this phenomenon. The most plausible explanation is the emergence of artifacts associated with high-norm tokens in the activation maps of larger models, which are used to store information (Darcet et al. 2023). Interestingly, this decrease in score is not perceptible in image perturbation-based XAI techniques, which reinforces the hypothesis that activation artifacts contribute to the reduced interpretability of explanations.
Steering XAI methods towards better alignment
We propose to use the PASTA-score to enhance the interpretability of an off-the-shelf XAI method, namely RISE (Petsiuk, Das, and Saenko 2018). Our approach is as follows: while RISE generates random masks and selects the one that has the best class scores $S_{proba}$, we introduce a regularization component based on the PASTA-score. Consequently, instead of rating masks using $S_{proba}$, we employ the following formula:
$w_{RISE+PASTA}=\lambda S_{PASTA}+(1-\lambda)S_{proba}\,$, (4)
where $\lambda\in[0,1]$ is a hyperparameter. When $\lambda=0$, the generated explanation aligns with the original RISE method. Conversely, if $\lambda=1$, the explanation produced corresponds to a scenario that maximizes the PASTA-score. Note that setting $\lambda=1$ would result in an explainer optimizing only for human preferences while neglecting the true behavior of the model, which may not yield useful explanations.
(a) $\lambda=0$
F = $0.0983$
P= $4.13$
(b) $\lambda=0.7$
F = $0.0874$
P = $4.19$
(c) $\lambda=1$
F = $0.0326$
P = $4.32$
(d) $\lambda=0$
F = $0.0692$
P = $2.82$
(e) $\lambda=0.7$
F = $0.0731$
P = $2.96$
(f) $\lambda=1$
F = $0.0674$
P = $3.15$
Figure 9: Optimized explanations derived through RISE adjusted with the PASTA-score. Label of the top images: home_or_hotel. Label of the bottom images: Renaissance. F denotes the faithfulness correlation score, P denotes the PASTA-score.
Upon analyzing the samples generated through the optimization process, we initially observe a slight improvement in the localization of highlighted objects. For instance, the explanation depicted in Figure 9(e) exhibits fewer indecisive zones and demonstrates enhanced precision compared to the explanation shown in Figure 9(d). Regarding the case where $\lambda=1$, we note that the explanation begins to hallucinate zones of interest while omitting others, like in Figure 9(c). Additionally, one can observe that the PASTA-score favors large heatmaps. However, the optimized explanations do not systematically overlap with the entire zone of the prediction, suggesting that alignment with the segmentation map of the object to assess the quality of saliency-based explanations, as conducted in previous studies (Karmani et al. 2024; Li et al. 2022b), may prove to be inadequate.
Conclusions
We introduce PASTA, a novel perceptual scoring method designed to benchmark XAI techniques in a human-centric manner. We collect a large-scale benchmark dataset (PASTA-dataset), and use it for an assessment of XAI explanations by human annotators. Based on this dataset, we develop an automated scoring method (PASTA-score) that spans previously unexplored modalities, effectively mimicking human preferences and allowing for circumventing resource-intensive user studies for applications that may benefit so. Deploying PASTA allows new quantitative observations: Our findings reveal a distinct preference for saliency-based explanations, identify a negative impact of backbone size, and demonstrate the potential to generate more human pleasant explanations without compromising faithfulness. These results not only align with human intuition but also corroborate visual examples, affirming the scalability and reliability of PASTA-score.
Limitations:
First, the PASTA-score is trained on specific datasets and explanation modalities, which may limit its generalizability to other unseen domains, especially those with domain-specific semantics. Second, the human preferences captured by the PASTA-dataset may inherit the intrinsic biases of human annotators. Third, although PASTA reduces the need for costly user studies, it remains an approximation of subjective human judgment and may overlook nuanced or task-specific interpretability needs, which may justify the need for more resource-intensive ad-hoc human interactions in downstream use cases.
Broader impact: